OpenStudy (anonymous):

On the same grid she will draw parallelogram PQRS similar to parallelogram ABCD with the ratio of corresponding side lengths PQ and AB equal to 2 : 1. If the coordinates of vertex P are (-4, 4), what could be the coordinates of vertex Q?

jimthompson5910 (jim_thompson5910):

is this a continuation of a prev problem?

jimthompson5910 (jim_thompson5910):

seems like something is missing, but I'm not sure

OpenStudy (anonymous):

no, nothing is missing. i'll post the answers for you

OpenStudy (anonymous):

(-8, 8) (-4, 10) (8, 3) (2, 4)

jimthompson5910 (jim_thompson5910):

is there a picture of PQRS ?

jimthompson5910 (jim_thompson5910):

or ABCD ?

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

AB is 3 units (since 5 - 2 = 3)

jimthompson5910 (jim_thompson5910):

the ratio of the side lengths of PQRS to ABCD is 2:1 So PQ is twice is large as AB So PQ = 2*AB PQ = 2*3 PQ = 6

jimthompson5910 (jim_thompson5910):

P is the point (-4, 4) So you need to find a point that is 6 units away from (-4, 4) to find Q

jimthompson5910 (jim_thompson5910):

The only point that fits is (2,4) since 2 - (-4) = 2+4 = 6

jimthompson5910 (jim_thompson5910):

So Q is (2,4)

OpenStudy (anonymous):

ohhhh' okayy. thank you (:

jimthompson5910 (jim_thompson5910):

you're welcome